OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

Use the equation y = mx +b where m is the slope, and plug in your values for x, y, m and solve for b. Then you'll have your equation! :)

OpenStudy (ddcamp):

Or use point slope form: y-y₁ = m(x-x₁) Where (x₁,y₁) is the point that you know.